-300P +1000 = 700(P-.05)

Answer by checkley71(8403)

-300P+1000=700(P-.05)
-300P+1000=700P-35
-300P-700P=-35-1000
-1000P=-1035
P=-1035/-1000
P=1.035 ANSWER.
PROOF.
-300*1.035+1000=700(1.035-.05)
-310.5+1000=700*.985
689.5=689.5